THe sum of two numbers is 4 less than twice the larger. If the larger is decreased by three times the smaller, the result is -20. Find the numbers.
Answer by josmiceli(19441)

a = larger
b = smaller
a + b = 2a - 4
subtract 2a from both sides
-a + b = -4
a - 3b = -20
add these equations
-2b = -24
b= 12
a - 3b = -20
a - 3*12 = -20
a - 36 = -20
a = 16
check
16 + 12 = 2*16 - 4
28 = 32 - 4
28 = 28
also
a - 3b = -20
16 - 3*12 = -20
16 - 36 = -20
-20 = -20
OK